Neutron stars are created when giant stars die in supernovas and their remaining cores

collapse to a state of immense density where protons and electrons combine to form

neutrons. A neutron star is ~1.4 times as massive as the sun and has radius of only

~10 km. For this neutron star compute its escape velocity. What percentage of the

speed of light does this correspond to? (Assume for the mass of the sun, M = 1.989 x 10^30 kg).
Answer #1

Escape Velocity=sqrt(2*G*Mass of neutron star/radius)=sqrt(2*6.67*10^(-11)*1.989*10^(30)*1.4/10000)=192734439.06059 m/s

Speed of light=299792458 m/s

% corresponding to speed of light=192734439.06059/299792458=64.2893%